Problem: The owner of a health foods store determines that the demand equation for selling a nutritional supplement is p=18-0.5√(x-4) where p is the price in dollars and x is the number of bottles demanded per week. How many bottles are demanded per week if the price of the supplement is $13.50?

Help So confused!!

Answer by stanbon(75887)

The owner of a health foods store determines that the demand equation for selling a nutritional supplement is p=18-0.5√(x-4) where p is the price in dollars and x is the number of bottles demanded per week. How many bottles are demanded per week if the price of the supplement is $13.50?
------------
1350 = 18-0.5√(x-4)
32/0.5 = sqrt(x-4)
64 = sqrt(x-4)
x-4 = 4096
x = 4100
==============
Cheers,
Stan H.